OpenStudy (anonymous):

how do you solve x^2-12x+4=0 ?

OpenStudy (saifoo.khan):

Quadratic formula!

OpenStudy (mertsj):

use the quadratic formula since it will not factor.

OpenStudy (anonymous):

or completing the square

OpenStudy (anonymous):

x^2-12 x+4 = 0 Subtract 4 from both sides: x^2-12 x = -4 Add 36 to both sides: x^2-12 x+36 = 32 Factor the left hand side: (x-6)^2 = 32 Take the square root of both sides: sqrt(x-6) = 4 sqrt(2) Eliminate the absolute value: x-6 = -4 sqrt(2) or x-6 = 4 sqrt(2) Add 6 to both sides: x = 6-4 sqrt(2) or x-6 = 4 sqrt(2) Add 6 to both sides: x = 6-4 sqrt(2) or x = 6+4 sqrt(2)

OpenStudy (anonymous):

^what she did :)